OpenStudy (anonymous):

the square root of 49a^4 Just wanted to make sure it is 7a^2 can someone check

jimthompson5910 (jim_thompson5910):

you are 100% correct, nice work

jimthompson5910 (jim_thompson5910):

sqrt(49) = 7 sqrt(a^4) = a^(4/2) = a^2

OpenStudy (anonymous):

thanks again! @jim_thompson5910

jimthompson5910 (jim_thompson5910):

yw